# Quantifying the Fermi paradox via passive SETI: a general framework

### v1

@@ 5.2 Limitations of this model
    Our analysis considers the existence of currently-observable EM signals within the galaxy, and assumes that all $r_{A}$ are equally likely. It may be the case, however, that the number of civilizations is time-dependent. This paper only considers the observation of signals last produced within about the past $10^{5}$ years, which is about five orders of magnitude less than the time during which the MW galaxy has existed. The distribution of observational windows, as well as stellar lifetimes and development times, is discussed in Kreifeldt (1971); Wallenhorst (1981). Future research may wish to apply these concepts to our model.
@@
    In this paper, we have relied upon the approximation $\delta\ll 1$, and we have taken $P=0.6\delta$. A more realistic treatment might re-derive Equation 3.2.1, for example, without assuming $P\ll 1$. Such a model would therefore consider signals emitted for of order 10,000 years or more.
@@ 6 Conclusion
In this paper, we have articulated a model which can provide a quantitative assessment of the Fermi paradox. This model has serious limitations, but we argue here that it can provide a generic framework in which to quantify the usefulness of SETI in respect to the existence of intelligent life. Our model shows that the lack of observation of life-produced EM signals places limitations on the number of EM-emitting civilizations. Given the presumptions in our model, we show that there is about a $99\%$ probability of observing at least one EM signal out of about $10^{2.8}$, if $\mathcal{N}=r_{*}f_{p}n_{e}f_{l}f_{i}f_{c}\approx 1$. Although this result is derived from a toy model, we have nonetheless shown that, at least in principle, one can rule out parts of $\mathcal{N}-l$ space merely from a null SETI result. As future research reduces the limitations discussed in Section 5.2, constraints on $\mathcal{N}-l$ space from the model presented in this paper may help provide insight into the resolution of the Fermi paradox.

### v2

@@ 4.3 Limitations of this model
    Our analysis considers the existence of currently-observable EM signals within the galaxy, and assumes that all signal radii are equally likely. It may be the case, however, that the number of civilizations is time-dependent. This paper only considers the observation of signals last produced within about the past $10^{5}$ years, which is about five orders of magnitude less than the time during which the MW galaxy has existed. The distribution of observational windows, as well as stellar lifetimes and development times, is discussed in Kreifeldt (1971); Wallenhorst (1981). Future research may wish to apply these concepts to our model.
@@
    In this paper, we have relied upon the approximation $\delta\ll 1$, and we have taken $P=0.6\delta$. A more realistic treatment might re-derive Equation 4.1, for example, without assuming $P\approx\delta\ll 1$. Such a model would therefore consider signals emitted for of order 10,000 years or more.
@@ 5 Conclusion
We have articulated a model which can provide a quantitative assessment of the Fermi paradox. This model has serious limitations, but we argue here that it can provide a generic framework in which to quantify the usefulness of SETI in respect to the existence of intelligent life. Our model shows that the lack of observation of life-produced EM signals places limitations on the number of EM-emitting civilizations. Given the presumptions in our model, we show that there is about a $99\%$ probability of observing at least one EM signal out of about $10^{2.8}$, if $\mathcal{N}=r_{*}f_{p}n_{e}f_{l}f_{i}f_{c}\approx 1$. Although this result is derived from a toy model, we have nonetheless shown that, at least in principle, one can rule out parts of $\mathcal{N}-l$ space merely from a null SETI result. As future research reduces the limitations discussed in Section 4.3, constraints on $\mathcal{N}-l$ space from the model presented in this paper may help provide insight into the resolution of the Fermi paradox.
Appendix A A Simple Two-Dimensional Geometrical Model
Here, we derive the previously-cited geometrical results associated with our geometrical model. This discussion will be generally similar to  Grimaldi (2017), except that we will provide an exact solution in the two-dimensional case, and we will show that in the simplest interpretation the location of Earth does not affect the likelihood of observation. The latter provides a very useful way of generalizing these results to more complex probability distributions. Our treatment differs further in that we then connect these results to the Drake equation, and argue that its parameter space can be constrained via a null result from SETI.
A.1 Introduction
These light fronts create an [^5] The annulus is non-concentric; for simplicity, we will simply refer to it as an “annulus”. annulus whose area may only be partly located within the MW galaxy. Let us discuss how to compute this. The location of the source is $(r_{0},\theta_{0})$, and Earth is denoted by the blue dot, of position $(r_{E},\theta_{E})$. Observation will correspond to an orientation where the Earth is in-between the two light-fronts. Hence, observation will only occur when
Figure 5: Here, we depict the basic geometry of our model. The MW galaxy has a radius “R”. The two light-fronts, emanating from an extraterrestrial source “S”, are labelled “A” and “B”. The source is a distance $r_{0}$ from the center of the MW galaxy, and $\tilde{r}_{A}$ is the distance from the origin to a point on light-front A. The distance of the Earth from the center of the Milky Way is labeled $r_{E}$”, and this figure depicts the condition defined by Equation 8.
For a single pair of light fronts, the probability of observation for a randomly-placed observer is the probability that this randomly-placed observer will fall in-between the light fronts; this is $\tilde{A}/\pi R^{2}$, where $\tilde{A}$ is the area that is both in-between the two light fronts and also within the MW galaxy. The area $\tilde{A}$ is circumscribed by the curves whose intersections with the edge of the MW galaxy are defined by the red and green dots. All points along the light-front AA ($B$) form angles $\tilde{\theta}_{A}$ ($\tilde{\theta}_{B}$), as shown in Figure 5. The distance from the center of the MW galaxy to any point along light-front AA ($B$) is $\tilde{r}_{A}$ ($\tilde{r}_{B}$). The red (green) dots are defined by $\tilde{r}_{A}=R$ ($\tilde{r}_{B}=R$), which implies that (keeping to just the red dots as an example)
$r_{A}^{2}=R^{2}+r_{0}^{2}-2Rr_{0}\cos(\theta_{0}-\tilde{\theta}_{A}),$ (9)
and thus
We can split useful $(r_{0},r_{A})$ parameter space into two parts: (1) where the argument of $\arccos$ is within [1,1][-1,1], and (2) where it is not. In the first case we will have two [^6] Except the singular case where $\Delta\theta=0$ when the argument is $1$. values for $\tilde{\theta}_{A}$, which we will call $\tilde{\theta}_{A1}=\theta_{0}-\Delta\theta$ and $\tilde{\theta}_{A2}=\theta_{0}+\Delta\theta$. This implies that
$R+r_{0}>r_{A}>R-r_{0}.$ (11)
A.2 Annulus area and probability of observation
In Figure 6, we depict just the essential components of our geometrical model sufficient to compute $\tilde{A}$. The inner area, $\tilde{A}_{A}$, can be bifurcated into two parts: the area along the inner concentric circle from one red dot to the other (shaded with red triangles), and the area along the edge of the MW (shaded with green x marks) from one red dot to the other. The latter is $R^{2}\left(\tilde{\theta}_{A2}-\tilde{\theta}_{A1}\right)/2=R^{2}\Delta\theta=\Delta\theta$, taking $R=1$; the former is $r_{A}^{2}\psi$, where $\psi$ is (as depicted in Figures 5 and 6) the angle between $r_{0}$ and $r_{A}$; $\psi$ appears to be
$\frac{\sin\psi}{\tilde{r}_{A}}$ $=\frac{\sin\Delta\theta}{r_{A}},$
$\Rightarrow\psi$ $=\sin^{-1}\left(\frac{\sqrt{4r_{0}^{2}-\left(1+r_{0}^{2}-r_{A}^{2}\right)^{2}}}{2r_{0}r_{A}}\right)\equiv\sin^{-1}\left(\chi\right),$
where we have taken $\tilde{r}_{A}=1$ so that $\psi$ is defined as in Figure 6. This angle isn’t quite right, however, since $\chi$ approaches 0 as $r_{A}$ approaches [^7] As $r_{A}$ approaches $1-r_{0}$, $1+r_{0}^{2}-r_{A}^{2}$ approaches $2r_{0}$; therefore the numerator in the definition of $\chi$ vanishes. $1-r_{0}$. As we can see from Figure 6, however, $\psi$ should be $\pi$ in this limit. Thus,
$\psi=\gamma\pi+\alpha\sin^{-1}\left(\chi\right),$ (12)
where $\alpha=-1$ ($\gamma=1$) when $\psi>\pi/2$ and +1+1 (0) otherwise. When these two regions meet at $\psi=\pi/2$,
$\sin^{-1}\left(\chi\right)=\pi-\sin^{-1}\left(\chi\right),$
which implies that $\chi=1$ and
$r_{A}^{2}+r_{0}^{2}+1-2r_{A}-2r_{0}+2r_{A}r_{0}=0.$
Equation 9 implies that $r_{A}=\sqrt{1+r_{0}^{2}-2r_{0}\cos\Delta\theta}$, which, when input into the equation just above, gives $r_{0}=\cos\Delta\theta$. Thus, $r_{A}=\sqrt{1+r_{0}^{2}-2r_{0}^{2}}=\sqrt{1-r_{0}^{2}}$ when [^8] We can see this also via the Pythagorean theorem, since the triangle in Figure 6 is a right triangle when $\psi=\pi/2$. $\psi=\pi/2$. Thus, $\alpha=-1$ ($\gamma=1$) when $r_{A}<\sqrt{1-r_{0}^{2}}$ and +1+1 (0) otherwise.
$\tilde{A}_{A}=\Delta\theta+r_{A}^{2}\psi-r_{0}\sin\Delta\theta.$ (13)
Therefore, the probability of a single observation is
[…]
Figure 6: Here, we depict the three areas that compose $\tilde{A}$: the area denoted by red triangles ($\psi r_{A}^{2}$), the area denoted by green x symbols ($\Delta\theta$), and their overlap. Thus, we subtract the last area from the sum of the first two. Note that here we take $R=1$.
Throughout this paper, we explore the mechanics of this approach via a toy model. This model is based upon several approximations, and, thus, the exact results are superfluous. Since for many practical purposes $r_{B}-r_{A}$ will be small compared to the radius of the MW, for example, we will in the next section implement this approximation and use it throughout the paper.
A.3 $\delta\equiv r_{B}-r_{A}\ll 1$ regime
Let us now approximate Equation 14 where $\delta\ll 1$. We begin with the bifurcation of the parameter space into whole source circles and partial source circles, as described in section A.1. When $\delta\equiv r_{B}-r_{A}\ll 1$, we may approximate the area $\tilde{A}$ as $\tilde{A}\approx\Omega\cdot\delta r_{A}$, where $\Omega=2\pi$ when the annulus is completely within the MW and $\Omega=2\psi$ otherwise. See Equation 12 for the definition of $\psi$. Thus,
$+2\pi\cdot\zeta[-rA+1-r0]\bigg),$ (15)
where
[…]
is the Heaviside step function. The expression $\zeta[rA-\left(1-r0\right)]\zeta[-rA+\left(1+r0\right)]=\zeta_{p}$ enforces the first conditional in Equation 14 (i.e., partial circles), and $\zeta[-rA+1-r0]=\zeta_{w}$ enforces the second in Equation 14 (i.e., whole circles). The average area over the MW is, therefore,
$\bar{A}_{\delta}$ $=\left(\int_{0}^{2}\int_{0}^{1}r_{0}\left(\zeta_{p}+\zeta_{w}\right)dr_{0}dr_{A}\right)^{-1}$
$\times\int_{0}^{2}\int_{0}^{1}r_{0}A_{\delta}dr_{0}dr_{A}$
$=0.6\pi\delta,$
$\Rightarrow P\approx P_{\delta}$ $=0.6\delta,$ (17)
A.4 The effect of the observer’s location
which is satisfied for all $r_{E}$. We can go further and use this to recompute Equation A.3. The probability of observation is the likelihood that an observer will fall in-between the light fronts $r_{A}$ and $r_{A}+\delta$, and, since the domain of $r_{A}$ is $0,1+r_{0}]$, this probability is $\delta/\left(1+r_{0}\right)$ if we assume that all $r_{A}$ are equally likely. Defining $\zeta=\zeta[-rA+\left(1+r_{0}\right)]$, the average probability for small $\delta$ is therefore
$P_{\delta}$ $=\left(\int_{0}^{2}\int_{0}^{1}r_{0}\zeta dr_{0}dr_{A}\right)^{-1}\times\int_{0}^{2}\int_{0}^{1}\frac{\delta r_{0}}{1+r_{0}}\zeta dr_{0}dr_{A}$
$=0.6\delta.$